OpenStudy (anonymous):

slope intercept form. I need help explaining step by step. (links in the comment)

OpenStudy (anonymous):

y = mx + b m is the slope. +1 on the x-axis means +m on the y-axis. b is the y-intercept. Where the line starts/When x = 0

OpenStudy (anonymous):

So, you want to try writing the equation? :P

OpenStudy (anonymous):

yes please. (:

OpenStudy (anonymous):

Did you understand what I just said??

OpenStudy (anonymous):

yes but how would you do it? I'm comparing new ideas with the techneuiqe i use/ (:

OpenStudy (anonymous):

b in this case is 5. When x = 0, y = 5. m is the slope. Rise/run, so I see after moving two units on the x-axis, we have gone DOWN 1 unit on the y-axis. Rise/run = -1/2 = -0.5 y = -0.5x+5 or y = -x/2+5

OpenStudy (anonymous):

Get it? o.O

OpenStudy (anonymous):

Oh crap, its this one!

OpenStudy (ledah):

slope intercept is pretty easy. want a tutorial on it..?

OpenStudy (anonymous):

If I'm reading it right, the equation for that one's the same as the other one D:

OpenStudy (ledah):

:o

OpenStudy (anonymous):

Oh nevermind, im silly >.< THANKS @aleroth